Students are given a list of ten vocabulary words to learn. In how many ways could four of the words be listed on a test?
Answer by fractalier(6550)

We say that is "ten chosen four at a time," and can be expressed as 10C4, or
(10!) / [(10-4)!(4!)] =
(10*9*8*7) / (4*3*2*1) =
210